OpenStudy (anonymous):

If |x+2|=|x-2|, then x=________ EXPLAIN PLEASE :)

OpenStudy (anonymous):

you really could just "guess and check"

OpenStudy (anonymous):

i need an explanation :(

OpenStudy (anonymous):

|...| means there is no negative so wich valu of x could do the same on both sides ?

OpenStudy (mathstudent55):

Let's say X is an expression in x and b is a nonnegative number. You solve |X| = b by solving X = b or -X = b

OpenStudy (anonymous):

so the answer is... 4? or -2?

OpenStudy (anonymous):

or zero?

OpenStudy (anonymous):

i mean, in my question though.

OpenStudy (anonymous):

my question is... How and Why? :) i just need an explanation :)

OpenStudy (mathstudent55):

|x+2|=|x-2| This is solved by turning it into these two equtions: x + 2 = |x - 2| or -(x + 2) = |x - 2| x + 2 = |x - 2| or -x - 2 = |x - 2| |x - 2| = x + 2 or |x - 2| = -x - 2 Each of the two equations turns into two equations: x - 2 = x + 2 or -(x - 2) = x + 2 or x - 2 = -x - 2 or -(x - 2) = -x - 2 no sol. or -x + 2 = x + 2 or 2x = 0 or -x + 2 = -x - 2 no sol. or -2x = 0 or x = 0 or no sol. no sol. or x = 0 or x = 0 or no sol. The only equations that lead to a solution give x = 0, so the only solution is x = 0.

OpenStudy (aravindg):

You can think of it graphically also |x+2|=|x-2| Simply means distance of x from -2 and 2 are equal .This is only possible when x lies on y axis ie x=0 .Done :)

OpenStudy (anonymous):

That answer is clearer. Thanks :D

OpenStudy (aravindg):

yw :)